OpenStudy (anonymous):

A bike has a depreciation value of $56 per year. If it was originally bought for $800, find the value of the bike after 4 years.

OpenStudy (anonymous):

\[$576\]

OpenStudy (triciaal):

Value, V = original cost - depreciation in $ per year for t number of years V = -56(t) + 800 after 4 years means t = 4 depreciation is 4(56) = 224 Value = 800 - 224 = 576

OpenStudy (anonymous):

thanks for explaining it!